# Near-Field Channel Estimation and Joint Angle-Range Recovery in XL-MIMO Systems: A Gridless Super-Resolution Approach

### v1

@@ Abstract
Existing near‑field channel estimation methods for extremely large‑scale MIMO (XL‑MIMO) typically discretize angle and range parameters jointly, resulting in large polar‑domain codebooks. This paper proposes a novel framework that formulates near‑field channel estimation as a gridless super‑resolution problem, eliminating the need for explicitly constructed codebooks. By employing a second‑order approximation of spherical‑wave steering vectors, the near-field channel is represented as a superposition of complex exponentials modulated by unknown waveforms. We demonstrate that these waveforms lie tightly in a common discrete chirp rate (DCR) subspace, with a dimension that scales as $\Theta(\sqrt{N})$ for an $N$-element array. By leveraging this structure and applying a lifting technique, we reformulate the non-convex problem as a convex program using regularized atomic norm minimization, which admits an equivalent semidefinite program. From the solution to the convex program, we obtain gridless angle estimates and derive closed‑form coarse range estimates, followed by refinement under the exact spherical model using gradient‑based nonlinear least squares. The proposed method avoids basis mismatch and exhaustive two‑dimensional grid searches while enabling accurate joint angle–range estimation with pilot budgets that scale sublinearly with array size in sparse multipath regimes. Simulations demonstrate accurate channel reconstruction and user localization across representative near‑field scenarios.
@@ I-B Notation and Organization of the Paper
Throughout the paper, we use lowercase (uppercase) bold characters to denote vectors (matrices). The $i$th element of a vector $\mathbf{x}$ is written as $\mathbf{x}]_{i}$. Similarly, the $(i,j)$th element of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}]_{i,j}$. For a vector $\mathbf{x}$, $\mathbf{x}^{T}$, $\overline{\mathbf{x}}$, $\mathbf{x}^{H}$, and $\|\mathbf{x}\|_{2}$ denote its transpose, complex conjugate, conjugate transpose, and $\ell_{2}$ norm, respectively. For a matrix $\mathbf{X}$, $\mathbf{X}^{T}$, $\overline{\mathbf{X}}$, $\mathbf{X}^{H}$, $\|\mathbf{X}\|_{F}$, $\operatorname{rank}(\mathbf{X})$, and $\operatorname{Tr}(\mathbf{X})$ denote the transpose, complex conjugate, conjugate transpose, Frobenius norm, rank, and trace of $\mathbf{X}$, respectively. The inner product of matrices $\mathbf{X}$ and $\mathbf{Y}$ is defined as $\langle\mathbf{X},\mathbf{Y}\rangle=\text{Tr}(\mathbf{Y}^{H}\mathbf{X})$ and $\langle\mathbf{X},\mathbf{Y}\rangle_{\mathbb{R}}=\text{Re}(\langle\mathbf{X},\mathbf{Y}\rangle)$. $\mathbf{X}\succeq 0$ indicates that $\mathbf{X}$ is positive semidefinite (PSD). We use $[N]$ and $\mathbf{I}_{N}$ to denote the set $\{0,1,\cdots,N-1\}$ and an $N\times N$ identity matrix, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the sets of real and complex numbers.
@@ III-B Discrete Chirp Rate Subspace
$\mathbf{g}(\phi_{l})\approx\mathbf{G}\mathbf{v}_{l},\penalty 10000\ l=1,\cdots,L,$ (16)
@@ III-C Channel Estimation via Atomic Norm Minimization
Similarly, defining $M=N_{\text{RF}}Q$ and $\mathcal{P}_{y}:\mathbb{C}^{P\times N}\rightarrow\mathbb{C}^{M}$ with $\mathcal{P}_{y}(\mathbf{X})]_{m}=\langle\mathbf{X},\sum_{n}[\mathbf{A}]_{m,n}\mathbf{q}_{n}\mathbf{e}^{H}_{n}\rangle$, we write the received pilot signal $\mathbf{y}$ compactly as
@@
$\text{minimize}_{\mathbf{X}}\quad\frac{1}{2}\|\mathbf{y}-\mathcal{P}_{y}(\mathbf{X})\|_{2}^{2}+\tau\|\mathbf{X}\|_{\mathcal{A}},$ (21)
@@ III-D Recovery Guarantees with Noisy Measurements
In this subsection, we establish the error bound for the regularized atomic norm minimization problem in (21), in the presence of additive Gaussian noise. To simplify the analysis, we assume that the noise follows the Gaussian distribution, i.e., $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{M})$, where $\sigma^{2}$ denotes the noise variance. Our analysis is inspired by the prior work on atomic norm denoising for line spectral estimation [Bhaskar-TSP13] and its extension to the multiple measurement vectors case [Li&Chi-TSP16].
@@
$\|\mathbf{Z}\|_{\mathcal{A}}^{\ast}=\sup_{\|\mathbf{X}\|_{\mathcal{A}}\leq 1}\langle\mathbf{Z},\mathbf{X}\rangle_{\mathbb{R}}=\sup_{\omega\in[0,1)}\|\mathbf{Z}\mathbf{a}(\omega)\|_{2}.$ (23)
@@
[…]
$\frac{1}{2\tau}\,\big\|\mathcal{P}_{y}(\widehat{\mathbf{X}}-\mathbf{X}_{h})\big\|_{2}^{2}\ \leq\ \|\mathbf{X}_{h}\|_{\mathcal{A}}.$ (24)
@@
[…]
Next, we show that when the noise is i.i.d. Gaussian, the expectation of the dual norm $\Big\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\,\Big\|_{\mathcal{A}}^{\ast}$ can be bounded.
Proposition 2.
Let $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{M})$. Define $q_{N,P}=\sqrt{N\log(4N)}+\sqrt{PN}$ and the operator norm $\|\mathcal{P}_{y}\|_{F\rightarrow 2}=\sup_{\|\mathbf{W}\|_{F}=1}\|\mathcal{P}_{y}(\mathbf{W})\|_{2}$. There exist constants $C,C^{\prime}>0$ such that
$\mathbb{E}\,\Big\|\mathcal{P}_{y}^{*}(\mathbf{n})\Big\|_{\mathcal{A}}^{*}\leq C\,\sigma\,\|\mathcal{P}_{y}\|_{F\to 2}\,q_{N,P},$ (25)
and for $\forall\,\delta\in(0,1)$,
$\mathbb{P}\!\left\{\Big\|\mathcal{P}_{y}^{*}(\mathbf{n})\Big\|_{\mathcal{A}}^{*}\leq C^{\prime}\,\sigma\,\|\mathcal{P}_{y}\|_{F\to 2}\,\Big(q_{N,P}+\sqrt{N\log(\frac{1}{\delta})}\Big)\right\}$ (26)
$\geq 1-\delta.$
The proof of Proposition 2 is provided in Appendix -C.
Using the results of (25) and (26) from Proposition 2, and by selecting an appropriate hyperparameter $\tau$, we have the following corollary.
Corollary 1.
Assume that $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{M})$.
  1. 1.
    If we take
    $\tau=C\,\sigma\,\|\mathcal{P}_{y}\|_{F\to 2}\,\Big(\sqrt{N\log(4N)}+\sqrt{PN}\Big),$ (27)
    the minimizer $\widehat{\mathbf{X}}$ of (21) satisfies
    $\mathbb{E}\,\big\|\mathcal{P}_{y}(\widehat{\mathbf{X}}-\mathbf{X}_{h})\big\|_{2}^{2}\ \leq\ 2\,\tau\,\|\mathbf{X}_{h}\|_{\mathcal{A}}.$ (28)
  2. 2.
    For any $\delta\in(0,1)$, if $\tau$ is chosen as
    $\begin{split}\tau=&C^{\prime}\,\sigma\,\|\mathcal{P}_{y}\|_{F\to 2}\,\times\\ &\Big(\sqrt{N\log(4N)}+\sqrt{PN}+\sqrt{N\log(\frac{1}{\delta})}\Big),\end{split}$ (29)
    then with probability at least $1-\delta$,
    $\big\|\mathcal{P}_{y}(\widehat{\mathbf{X}}-\mathbf{X}_{h})\big\|_{2}^{2}\ \leq\ 2\,\tau\,\|\mathbf{X}_{h}\|_{\mathcal{A}}.$ (30)
Proof:
By selecting $\tau$ according to (27), Proposition 2 implies that $\mathbb{E}\big\|\tfrac{1}{\tau}\mathcal{P}_{y}^{*}(\mathbf{n})\big\|_{\mathcal{A}}^{*}\leq 1$. Taking expectations in (24) with respect to the randomness of the noise yields (28). Furthermore, by choosing $\tau$ as in (29), Proposition 2 ensures $\big\|\tfrac{1}{\tau}\mathcal{P}_{y}^{*}(\mathbf{n})\big\|_{\mathcal{A}}^{*}\leq 1$ with probability at least $1-\delta$. Thus, (30) holds with probability at least $1-\delta$. ∎
@@ IV Joint Angle and Range Estimation
This section develops a two-stage strategy for joint recovery of angle and range parameters from the solution to the convex problem (22), i.e., $\{\widehat{\mathbf{X}},\widehat{\mathbf{u}},\widehat{\mathbf{V}}\}$. Stage-I extracts gridless angle and coarse range estimates directly from $\{\widehat{\mathbf{X}},\widehat{\mathbf{u}},\widehat{\mathbf{V}}\}$. Stage-II refines both angle and range estimates under the exact spherical-wave steering model via gradient-based nonlinear least squares.
@@ IV-A Gridless Angle and Coarse Range Estimation
An estimate of the vector $\alpha_{1}\mathbf{v}_{1}^{T},\cdots,\alpha_{L}\mathbf{v}_{L}^{T}]^{T}$ is obtained via least squares. Note that we cannot resolve the intrinsic scaling ambiguity between each $\alpha_{l}$ and the corresponding $\mathbf{v}_{l}$ due to their multiplications.
@@ IV-C Computational Complexity
In this subsection, we analyze the computational complexity of the proposed method in Algorithm 1. Since Stage-I requires solving the SDP in (22), its cost dominates that of Stage-II. Therefore, we focus solely on the complexity of Stage-I. According to [Boyd-SDP-SIAM], when an SDP is solved using interior-point methods (e.g., CVX/SDPT3), the computational complexity is $\mathcal{O}((N_{1})^{2}(N_{2})^{2.5})$, where $N_{1}$ is the number of optimization variables and $N_{2}\times N_{2}$ is the size of the PSD matrix. Therefore, the computational complexity for solving (22) is $\mathcal{O}((NP)^{2}(N+P)^{2.5})$. When first‑order solvers (e.g., ADMM) are applied to solve the SDP, the computational complexity can be reduced to $\mathcal{O}(K_{\rm 1st}(N+P)^{3})$, where $K_{\rm 1st}$ is the number of first-order iterations. Therefore, first-order methods significantly reduce the computational complexity for solving (22).
@@ V Numerical Simulations
In this section, we present simulation results to validate the effectiveness of the proposed near-field channel estimation scheme. To evaluate the contribution of each stage in Algorithm 1, we consider two variants of our approach: (i) a Stage-I-only method, denoted as “ANM(SI)”, and (ii) a two-stage method that employs both Stage-I and Stage-II steps, denoted as “ANM(SI+SII)”. We compare these methods with four representative baselines: the on-grid angular-domain SW-OMP [Rodriguez-2018TWC], the on-grid polar-domain P-SOMP [Cui-ToC2022], the off-grid angular-domain SS-SIGW-OLS [Gonzalez-TWC2021], and the off-grid polar-domain P-SIGW [Cui-ToC2022].
The simulations consider a MIMO communication system in which the BS is equipped with $N=256$ antennas, operates at a carrier frequency of $f_{c}=100\text{GHz}$, and uses $N_{\text{RF}}=4$ RF chains. The corresponding Rayleigh distance for this configuration is approximately $98\text{m}$. The user transmits a pilot sequence of length $Q=32$. The near-field channel is assumed to consist of $L=4$ propagation paths, including one LOS path and three NLOS paths. The dimension of the DCR subspace is set to $P=10$. The regularization hyperparameter $\tau$ is selected according to (27), where $C$ is optimized at different SNR levels. Unless otherwise specified, the simulations use the parameter configuration described above.
@@ V-A Selection of DCR Subspace Dimension
As shown in Fig. 2(a), selecting $P$ around $9$ or $10$ yields the lowest NMSE across SNR levels of 10, 20, and 30 dB. This indicates that a moderate subspace dimension sufficiently captures the near-field structure while avoiding excessive model complexity. Fig. 2(b) shows that the angle estimation RMSE reaches its minimum near $P=10$ at moderate and high SNRs, and near $P=7$ at the SNR level of 10 dB. Similar trends are observed for range estimation in Fig. 2(c), where comparable $P$ values achieve the best RMSE. These results demonstrate that an appropriate DCR subspace dimension effectively characterizes the spherical wavefront curvature and enables reliable angle and range recovery. In all subsequent simulations, we set $P=10$, which achieves an effective balance between estimation accuracy and computational efficiency.

### v2

@@ Abstract
Existing near‑field channel estimation methods for extremely large‑scale MIMO (XL‑MIMO) typically discretize angle and range parameters jointly, resulting in large polar‑domain codebooks. This paper proposes a novel framework that formulates near‑field channel estimation as a gridless super‑resolution problem, eliminating the need for explicitly constructed codebooks. By employing a second‑order approximation of spherical‑wave steering vectors, the near-field channel is represented as a superposition of complex exponentials modulated by unknown waveforms. We demonstrate that these waveforms lie tightly in a common discrete chirp rate (DCR) subspace, with a dimension that scales as $\Theta(\sqrt{N})$ for an $N$-element array. By leveraging this structure and applying a lifting technique, we reformulate the non-convex problem as a convex program using regularized atomic norm minimization, which admits an equivalent semidefinite program. From the solution to the convex program, we obtain gridless angle estimates and derive closed‑form coarse range estimates, followed by refinement under the exact spherical model using gradient‑based nonlinear least squares. The proposed method avoids basis mismatch and exhaustive two‑dimensional grid searches while enabling accurate joint angle-range estimation with pilot budgets that scale sublinearly with array size in sparse multipath regimes. Simulations demonstrate accurate channel reconstruction and user localization across representative near‑field scenarios.
@@ I-B Notation and Organization of the Paper
Throughout the paper, we use lowercase (uppercase) bold characters to denote vectors (matrices). The $i$th element of a vector $\mathbf{x}$ is written as $\mathbf{x}]_{i}$. Similarly, the $(i,j)$th element of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}]_{i,j}$. For a vector $\mathbf{x}$, $\mathbf{x}^{T}$, $\overline{\mathbf{x}}$, $\mathbf{x}^{H}$, and $\|\mathbf{x}\|_{2}$ denote its transpose, complex conjugate, conjugate transpose, and $\ell_{2}$ norm, respectively. For a matrix $\mathbf{X}$, $\mathbf{X}^{T}$, $\overline{\mathbf{X}}$, $\mathbf{X}^{H}$, $\|\mathbf{X}\|_{2}$, $\|\mathbf{X}\|_{F}$, $\operatorname{rank}(\mathbf{X})$, and $\operatorname{Tr}(\mathbf{X})$ denote the transpose, complex conjugate, conjugate transpose, operator norm, Frobenius norm, rank, and trace of $\mathbf{X}$, respectively. The inner product of matrices $\mathbf{X}$ and $\mathbf{Y}$ is defined as $\langle\mathbf{X},\mathbf{Y}\rangle=\text{Tr}(\mathbf{Y}^{H}\mathbf{X})$ and $\langle\mathbf{X},\mathbf{Y}\rangle_{\mathbb{R}}=\text{Re}(\langle\mathbf{X},\mathbf{Y}\rangle)$. $\mathbf{X}\succeq 0$ indicates that $\mathbf{X}$ is positive semidefinite (PSD). We use $[N]$ and $\mathbf{I}_{N}$ to denote the set $\{0,1,\cdots,N-1\}$ and an $N\times N$ identity matrix, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the sets of real and complex numbers.
@@ III-B Discrete Chirp Rate Subspace
$\mathbf{g}(\phi_{l})\approx\mathbf{G}\mathbf{v}_{l},~l=1,\cdots,L,$ (16)
@@ III-C Channel Estimation via Atomic Norm Minimization
Similarly, defining $M=N_{\text{RF}}Q$ and $\mathcal{P}_{y}:\mathbb{C}^{P\times N}\rightarrow\mathbb{C}^{M}$ with $\mathcal{P}_{y}(\mathbf{X})]_{m}=\sum_{n}[\mathbf{A}]_{m,n}\langle\mathbf{X},\mathbf{q}_{n}\mathbf{e}^{H}_{n}\rangle$, we write the received pilot signal $\mathbf{y}$ compactly as
@@
$\underset{\mathbf{X}}{\text{minimize}}\quad\frac{1}{2}\|\mathbf{y}-\mathcal{P}_{y}(\mathbf{X})\|_{2}^{2}+\tau\|\mathbf{X}\|_{\mathcal{A}},$ (21)
@@ III-D Recovery Guarantees with Noisy Measurements
In this subsection, we establish the error bound for the regularized atomic norm minimization problem in (21), in the presence of additive Gaussian noise. To simplify the analysis, we assume that the noise follows the Gaussian distribution, i.e., $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{M})$, where $\sigma^{2}$ denotes the noise variance. Our analysis is inspired by the prior work on atomic norm denoising for line spectral estimation [Bhaskar-TSP13].
@@
$\|\mathbf{Z}\|_{\mathcal{A}}^{\ast}=\sup_{\|\mathbf{X}\|_{\mathcal{A}}\leq 1}\langle\mathbf{Z},\mathbf{X}\rangle_{\mathbb{R}}=\sup_{\omega\in[0,1)}\|\mathbf{Z}\overline{\mathbf{a}(\omega)}\|_{2}.$ (23)
@@
Let $\mathcal{P}^{\ast}_{y}:\mathbb{C}^{M}\mapsto\mathbb{C}^{P\times N}$ be the adjoint operator of $\mathcal{P}_{y}$. If $\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast}\leq\ \tau$, then every minimizer $\widehat{\mathbf{X}}$ of (21) satisfies
$\frac{1}{N}\big\|\mathcal{P}_{y}(\widehat{\mathbf{X}}-\mathbf{X}_{h})\big\|_{2}^{2}\ \leq\ \frac{2\tau}{N}\|\mathbf{X}_{h}\|_{\mathcal{A}}.$ (24)
@@
Theorem 1 shows that when the regularization hyperparameter $\tau$ is appropriately selected such that $\tau\geq\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast}$, the recovery mean squared error of the low-rank matrix $\mathbf{X}_{h}$ from the SDP solution $\widehat{\mathbf{X}}$, when projected onto the observed entries, i.e., $\frac{1}{N}\|\mathcal{P}_{y}(\widehat{\mathbf{X}}-\mathbf{X}_{h})\|_{2}^{2}$, is upper bounded by $\frac{2\tau}{N}$ times the atomic norm of $\mathbf{X}_{h}$, which is typically small due to the low-rankness of $\mathbf{X}_{h}$.
Next, we show that when the noise is i.i.d. Gaussian, the expectation of the dual norm, i.e., $\mathbb{E}(\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast})$, can be bounded.
Theorem 2.
Assume that the entries of $\mathbf{n}$ are i.i.d. Gaussian and $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{M})$. Let $\tau=\sqrt{1+\frac{1}{\log N}}\sqrt{\log N+\log(4\pi\log N)+1}\|\mathbf{A}\|_{2}\|\mathbf{G}\|_{F}\sigma$. Then we have
$\mathbb{E}(\|\mathcal{P}_{y}^{\ast}(\mathbf{n})\|_{\mathcal{A}}^{\ast})\leq\tau.$ (25)
Furthermore, the minimizer $\widehat{\mathbf{X}}$ of (21) satisfies
$\frac{1}{N}\mathbb{E}\,\big\|\mathcal{P}_{y}(\widehat{\mathbf{X}}-\mathbf{X}_{h})\big\|_{2}^{2}\ \leq\ \frac{2\,\tau\,}{N}\|\mathbf{X}_{h}\|_{\mathcal{A}}.$ (26)
The proof of the key result in (25) above is provided in Appendix -C. The recovery mean squared error bound in (26) is obtained immediately by applying Theorem 1 and (25). Note that the error bound (26) is as tight as the standard error bound results derived in the atomic norm denoising for line spectral estimation problem [Bhaskar-TSP13, Li&Chi-TSP16].
@@ IV Joint Angle and Range Estimation
This section develops a two-stage strategy for joint recovery of angle and range parameters from the solution to the convex problem (22), i.e., $\{\widehat{\mathbf{X}},\widehat{\mathbf{u}},\widehat{\mathbf{V}}\}$. Stage-I extracts gridless angle and coarse range estimates directly from $\{\widehat{\mathbf{X}},\widehat{\mathbf{u}},\widehat{\mathbf{V}}\}$. Stage-II refines both angle and range estimates under the exact spherical-wave steering model via gradient-based methods.
@@ IV-A Gridless Angle and Coarse Range Estimation
An estimate of the vector $\alpha_{1}\mathbf{v}_{1}^{T},\cdots,\alpha_{L}\mathbf{v}_{L}^{T}]^{T}$ is obtained via least squares. Note that we cannot resolve the intrinsic scaling ambiguity between each $\alpha_{l}$ and the corresponding $\mathbf{v}_{l}$ due to their multiplication.
@@ IV-C Computational Complexity
In this subsection, we analyze the computational complexity of the proposed method in Algorithm 1. Since Stage-I requires solving the SDP in (22), its cost dominates that of Stage-II. Therefore, we focus solely on the complexity of Stage-I. According to [Boyd-SDP-SIAM], when an SDP is solved using interior-point methods (e.g., CVX/SDPT3), the computational complexity is $\mathcal{O}((N_{1})^{2}(N_{2})^{2.5})$, where $N_{1}$ is the number of optimization variables and $N_{2}\times N_{2}$ is the size of the PSD matrix. Therefore, the computational complexity for solving (22) is $\mathcal{O}((NP)^{2}(N+P)^{2.5})$. When first‑order solvers (e.g., ADMM) are applied to solve the SDP, the computational complexity can be reduced to $\mathcal{O}(K_{\rm 1st}(N+P)^{3})$, where $K_{\rm 1st}$ is the number of first-order iterations. Therefore, first-order methods significantly reduce the computational complexity for solving (22). We leave the development of fast solvers, such as ADMM, to future work.
@@ V Numerical Simulations
In this section, we present simulation results to validate the effectiveness of the proposed near-field channel estimation scheme. To evaluate the contribution of each stage in Algorithm 1, we consider two variants of our approach: (i) a Stage-I-only method, denoted as “ANM(SI)”, and (ii) a two-stage method that employs both Stage-I and Stage-II steps, denoted as “ANM(SI+SII)”. We compare these methods with four representative baselines: the on-grid angular-domain SW-OMP [Rodriguez-2018TWC], the on-grid polar-domain P-SOMP [^1] Note that the original source code of the P-SOMP algorithm sets the number of propagation paths parameter to $2L$ (i.e., double the actual number of paths). For consistent comparisons, we set this parameter to $L$ in our simulations. [Cui-ToC2022], the off-grid angular-domain SS-SIGW-OLS [Gonzalez-TWC2021], and the off-grid polar-domain P-SIGW [Cui-ToC2022].
The simulations consider a MIMO communication system in which the BS is equipped with $N=256$ antennas, operates at a carrier frequency of $f_{c}=100\text{GHz}$, and uses $N_{\text{RF}}=4$ RF chains. The corresponding Rayleigh distance for this configuration is approximately $98\text{m}$. The user transmits a pilot sequence of length $Q=32$. The near-field channel is assumed to consist of $L=4$ propagation paths, including one LOS path and three NLOS paths. The dimension of the DCR subspace is set to $P=10$. The regularization hyperparameter $\tau$ is selected according to Theorem 2. Unless otherwise specified, the simulations use the parameter configuration described above.
@@ V-A Selection of DCR Subspace Dimension
As shown in Fig. 2(a), selecting $P$ around $9$ or $10$ yields the lowest NMSE across SNR levels of 10, 20, and 30 dB. This indicates that a moderate subspace dimension sufficiently captures the near-field structure while avoiding excessive model complexity. Fig. 2(b) shows that the angle estimation RMSE reaches its minimum near $P=10$ at moderate and high SNRs, and near $P=7$ at the SNR level of 10 dB. Similar trends are observed for range estimation in Fig. 2(c), where comparable $P$ values achieve the best RMSE. These results demonstrate that an appropriate DCR subspace dimension effectively characterizes the spherical wavefront curvature and enables reliable angle and range recovery for our proposed approach. In all subsequent simulations, we set $P=10$, which achieves an effective balance between estimation accuracy and computational efficiency.